# A Survey on Diffusion Models for Recommender Systems

### v1

@@ 3.2.1. Diffusion for Collaborative Recommendation
Instead of manipulating over the entire interaction matrix, other works propose to employ the diffusion-denoising process over the single user-level interaction records with parameterized learnable functions (Walker et al., 2022; Wang et al., 2023b; Bénédict et al., 2023; Jiangzhou et al., 2024; Zhu et al., 2024). DiffRec (Wang et al., 2023b) takes as input each user’s binary interaction vector, i.e., $\mathbf{x}_{0}=\{0,1\}^{|\mathcal{I}|}$ with the $i$-th binary element implies whether the user has interacted with item $i$ or not, and adopts the classical DDPM framework to enable the generative recommendation paradigm. Two improvements over DDPM are proposed to ensure better personalized recommendations. (1) The authors reduce the noise scales and forward diffusion steps to control the corruption over user interaction records. (2) DiffRec is optimized by directly predicting the target interaction history $\mathbf{x}_{0}$ instead of the noise $\epsilon$ to be eliminated at each step, which is more intuitive and empirically training-stable. Moreover, DiffRec is further extended into two versions. (1) L-DiffRec compresses the input vector via item clustering and thereby enable latent diffusion to reduce the resource costs for large-scale item prediction. (2) T-DiffRec incorporates the temporal information by applying positional reweighting to the input interaction vector $\mathbf{x}_{0}$ to further capture the temporal dynamics of user preferences. It is worth noting involving the temporal information actually leads to the field of sequence recommendation (i.e., context-aware recommendation in Section 3.2.2), but the major designs and contributions of DiffRec (Wang et al., 2023b) are still within the scope of collaborative recommendation. Other works in this type generally follow such a setting. For instance, CODIGEM (Walker et al., 2022) utilizes multiple autoencoders (AEs) to model the reverse denoising generation yet only leverages the first AE for interaction prediction during the inference phase. Bénédict et al. (2023) propose 1D binomial diffusion to explicitly model the binary user-level interaction vectors with a Bernoulli process. GiffCF (Zhu et al., 2024) converts the user-level interaction vector into an item-item similarity graph, and employs a smoothed graph-based diffusion-denoising process using the heat equations, which leverages the advantages of both diffusion models and graph signal processing.
@@ 3.2.2. Diffusion for Context-aware Recommendation
First and foremost, a number of works aim to involve temporal/sequential information for diffusion-based context-aware recommender models (Yang et al., 2024c; Li et al., 2023a; Du et al., 2023; Wang et al., 2024c; Niu et al., 2024; Wang et al., 2024d; Li et al., 2024c). These studies intend to conduct conditional diffusion-based generative recommendation by taking the representations of user behavior sequences as the guidance for reverse denoising process. For example, DreamRec (Yang et al., 2024c) uses a Transformer (Vaswani et al., 2017) encoder to create guidance representations from user behavior sequences and employs a diffusion model to explore the underlying distribution of item space, generating an oracle next-item embedding that aligns with user preferences without the need for negative sampling. As a follow-up work, DiffRIS (Niu et al., 2024) improves the guidance representations from behavior sequences by explicitly modeling the long- and short-term user interests via delicately designed multi-scale CNN and residual LSTM modules. DimeRec (Li et al., 2024c), on the other hand, strives to optimize the diffusion-denoising process. To be specific, for each training sample, the model introduces noise to the target item embedding using geodesic random walk (De Bortoli et al., 2022) on a spherical space, ensuring isotropic and small noise insertion that approximates Gaussian distribution in the tangent space. The reverse denoising process is conducted under the guidance of user sequence representation, and is jointly optimized with the objective of traditional sequential recommender in a multi-task manner. Moreover, instead of integrating the sequential behaviors into one compact representation beforehand for guidance, other works (Du et al., 2023; Li et al., 2023a; Wang et al., 2024c) attempt to directly incorporate a sequence of historical item embeddings during the reverse generation process, resulting in implicit conditions on the sequential information.
@@ 3.2.3. Diffusion for Other Applications
While collaborative and context-aware diffusion-based recommender models focus on precisely estimating the user preference towards a target item given the in-domain recommendation data (i.e., interaction records and contextual information), diffusion models are also adapted as the core generative functions for other online application tasks. These tasks are closely related to the recommendation scenarios but somehow differ from the aforementioned two categories (i.e., collaborative or context-aware recommendation) in problem formulations and solution paradigms. Hence, we generally categorize them as other applications and briefly introduce these research works as follows.
@@
DMSR (Tomasi et al., 2024) attempts to optimize the slate recommendation with the help of diffusion models. Different from traditional recommender models that usually estimate the user preference towards each individual item, the slate recommendation aims to optimize the entire collection of items (i.e., a slate/group/bundle) presented to the user at once. Hence, the model should comprehensively consider the in-slate mutual effects (e.g., diversity) and overall utilities, thereby suffering from the complex combinatorial choice space. To this end, diffusion models turn out a promising solution to directly generate a set of latent vectors to construct the slate via item mapping. DMSR adopts the diffusion transformer architecture (Peebles and Xie, 2022), and conducts the reverse denoising process over the corrupted in-slate item representations under the guidance of contextual information like user preferences or textual queries. The denoised latent vectors are then decoded back into the discrete item space, resulting in a slate recommendation that maximizes user satisfaction by balancing the relevance and diversity.
@@ 3.2.4. Discussion
Due to the exceptional generative capabilities and flexibility with different backbones and downstream tasks, diffusion-based recommender models have turned out one of the the state-of-the-art generative recommendation paradigms by learning the underlying distribution of user preferences and item characteristics. From the aforementioned discussion, we can clearly observe the potential advantages through the adaptation of diffusion models as recommenders, including but not limited to capturing the complex and non-linear relationships between users and items and therefore generating diverse and novel recommendations. In this section, as shown in Figure 5, we would like to further discuss this line of research from the following three key perspectives when constructing the diffusion-based recommender models: (1) what to diffuse, (2) what is the guidance, and (3) how to accelerate.
@@
What to Diffuse? As shown in Figure 5, there are generally two types of inputs for the diffusion process: discrete user preference distribution and continuous latent representations. The first choice is to diffuse the interaction matrix or the probabilistic distribution of user preferences over the entire item space (Bénédict et al., 2023; Wang et al., 2023b), e.g., $\mathbf{x}_{0}=\{0,1\}^{|\mathcal{I}|}$. Although such kind of methods is simple and straightforward to directly denoise and generate the target user preference distribution towards potentially unobserved interactions, it is not scalable when facing a large volume of items due to the tremendously large input space and extensive resource requirements. Moreover, other contextual information like multi-modal knowledge can only be involved in the guidance part, which limits the model design. As a result, the second choice proves a more coherent, scalable and commonly adopted solution to conduct the diffusion-denoising process in the continuous latent space (Yu et al., 2023a; Wang et al., 2024e; Li et al., 2024f; He et al., 2024a). In this way, the diffusion-based recommenders are able to first integrate multi-source knowledge (e.g., spatial-temporal data and social networks in addition to collaborative signals) into a unified representation via specially designed encoders, and employ the diffusion model to learn the underlying latent distributions, which is more compatible with various recommendation tasks like POI recommendation or social recommendation. The denoised latent representations, usually serving as the user preference vectors or target item representations, are then used to retrieve the top-$K$ relevant items through vector matching for the final ranked item list.
@@
What is the Guidance (Optional)? As depicted in Figure 5, the guidance serves as the conditioning information to tailor the reverse denoising process for either discrete user preference distributions or continuous latent representations. It plays a crucial role in ensuring that the generated recommendations are not only high-quality but also relevant and personalized for each user. Although optional, the guidance mechanism has been widely adopted in diffusion-based recommender models, especially for context-aware recommendation where various context information can be fused into the conditioning signals (Yang et al., 2024c; Niu et al., 2024; Li et al., 2024c).
@@
How to Accelerate? By determining the aforementioned two key factors (i.e., contents to be diffused and the optional guidance), one can already construct the overall architecture of diffusion-based recommender models. However, the inference efficiency is another important topic when directly adapting diffusion models as recommenders, since online recommender systems are usually real-time services and extremely time-sensitive, where the user request should be responded within around tens of milliseconds (Lin et al., 2024b, 2023a). Diffusion models can be computationally expensive due to the multi-step denoising process, especially when dealing with large-scale recommender systems. Several strategies have been proposed to balance the trade-off between recommendation quality and computational efficiency. For example, Lin et al. (2024a) and Wang et al. (2023b) propose to start the denoising inference from a meaningful input (e.g., outputs from the previous stages or partially corrupted user preference representations) instead of the pure Gaussian noise in traditional diffusion models, which does not only avoid totally eliminating the important personalized information, but also accelerate the inference phase with much fewer denoising steps. Other works also attempt to speed up the denoising process by adopting the early-stopping strategy (Lin et al., 2024a) or one-step inference (Yu et al., 2023a). Despite these preliminary explorations, there still lack the systematic studies over the general acceleration strategies for the inference of diffusion-based recommender models, formulating a promising direction for future research over which we would cast further discussions in Section 4.
@@ 3.3. Diffusion for Content Presentation
The two stages discussed above (i.e., data engineering & encoding, and recommender models) concentrate on selecting and arranging the optimal ranked item list for recommendation. In this section, the phase of content presentation focuses on customizing and seeking the best presentation strategy of recommended items for different users (e.g., individualized titles or thumbnails). This is non-trivial to not only consider the user’s potential preferences based the historical and contextual information but also involve the visual appearance correlation and coherence. While earlier attempts for content presentation heavily rely on the manual design and pre-defined page arrangement, the emerging of top-tier generative models, especially diffusion models like Stable Diffusion (Rombach et al., 2022b), points out a promising direction to automate and even personalize the content generation for the item display. According to the different types of contents to be generated by diffusion models, we divide the existing research works in this line into three categories: (1) fashion generation, (2) ad creative generation, and (3) general content generation.
@@ 3.3.2. Ad Creative Generation
Compared with fashion outfits, ad creative generation covers a broader range of visual contents, including but not limited to clothing, cosmetics, and other promotional images. Previously, the creatives of ad campaigns are designed with great care and manual efforts. However, this human-designer-centered approach are non-scalable and suboptimal, because the designer often struggles to fully consider the global preferences of targeted user groups, and also fails to adapt to different online advertising scenarios. A common solution is to show the original product image with additional design elements adaptively generated by some neural models, e.g., language models to customize the headlines and captions (Mita et al., 2023; Kanungo et al., 2021). Moreover, since the ad creative aims to promote the sale or click rate of a specific product, there are usually more constraints for automatic ad creative generation. For example, we have to maintain the original target product within the generated or edited creative image. Besides, the seller and designer might have additional guidelines and instructions when employing artificial intelligence generated content (AIGC) for creative modification or generation, e.g., the overall visual styles and color themes should be coherent and consistent. To this end, diffusion models are becoming increasingly popular to facilitate the ad creative generation with inpainting/outpainting mode (Lugmayr et al., 2022) and text-to-image controllable generation (Abdollahpouri et al., 2017).
Shilova et al. (2023) propose the generative creative optimization task, and utilize user interest signals to personalize the ad creative generation with the outpainting mode of stable diffusion (Rombach et al., 2022b). The proposed AdBooster first masks out the background of the original product image (e.g., a model with certain fashion outfits), and then finetunes a stable diffusion model to outpaints the masked background with textual prompt guidance based on the user query and contextual information.
@@
More recently, Czapp et al. (2024) employ diffusion models for the creation of eye-catching personalized product images to increase user engagement with recommendations in online retargeting campaigns for real-world industrial applications. Although still adopting the basic fill-in-the-background paradigm for ad creative generation, the authors improve the overall pipeline from the following aspects. Firstly, the position and size of the masked product object would be further adjusted to the center of the image according to the placement for emphasis. Secondly, an optional edge detection is introduced to further reinforce the contours with the background mask. Finally, the authors use LinUCB (Li et al., 2010; Chu et al., 2011) contextual multi-armed bandit algorithm to select the prompt with the highest predicted CTR from a pre-defined pool of prompts associated with the product category in the given context (defined by user, item, and location features), which is designed to balance the generative personalization and online latency constraint.
@@ 3.3.3. General Content Generation
The exploratory research works (Mukande et al., 2024; Shen et al., 2024; Wang et al., 2023a) generally follow the similar framework, where the diffusion-based content generator is guided by the conditions extracted by a hybrid instructor module. The content to be generated by diffusion models is no longer limited to images of clothing or ad creatives, but has already extended to a broader range of modalities. The hybrid instructor module usually involves large language models (LLMs) to deepen the understanding of user intents assisted by other traditional models like sequential models and graph-enhanced models. Built upon this basic setting, Wang et al. (2023a) propose GereRec to further refine the formulation and deepen the integration of the diffusion-based AIGC module with the overall recommendation pipeline. GereRec not only generates the personalized content, but also supports a variety of operations to further adjust the item content to maximize the user-oriented utilities, e.g., thumbnail generation and selection, caption customization, and domain-specific fidelity checks. Although the research in this line are still preliminary and not well validated on large-scale platforms, we believe this is a direction worth delving into, which could potentially lead to the next-generation generative recommender paradigms.

### v2

@@ 3.2.1. Diffusion for Collaborative Recommendation
Instead of manipulating over the entire interaction matrix, other works propose to employ the diffusion-denoising process over the single user-level interaction records with parameterized learnable functions (Walker et al., 2022; Wang et al., 2023b; Bénédict et al., 2023; Jiangzhou et al., 2024; Zhu et al., 2024). DiffRec (Wang et al., 2023b) takes as input each user’s binary interaction vector, i.e., $\mathbf{x}_{0}=\{0,1\}^{|\mathcal{I}|}$ with the $i$-th binary element implies whether the user has interacted with item $i$ or not, and adopts the classical DDPM framework to enable the generative recommendation paradigm. Two improvements over DDPM are proposed to ensure better personalized recommendations. (1) The authors reduce the noise scales and forward diffusion steps to control the corruption over user interaction records. (2) DiffRec is optimized by directly predicting the target interaction history $\mathbf{x}_{0}$ instead of the noise $\epsilon$ to be eliminated at each step, which is more intuitive and empirically training-stable. Moreover, DiffRec is further extended into two versions. (1) L-DiffRec compresses the input vector via item clustering and thereby enables latent diffusion to reduce the resource costs for large-scale item prediction. (2) T-DiffRec incorporates the temporal information by applying positional reweighting to the input interaction vector $\mathbf{x}_{0}$ to further capture the temporal dynamics of user preferences. It is worth noting involving the temporal information actually leads to the field of sequence recommendation (i.e., context-aware recommendation in Section 3.2.2), but the major designs and contributions of DiffRec (Wang et al., 2023b) are still within the scope of collaborative recommendation. Other works of this type generally follow such a setting. For instance, CODIGEM (Walker et al., 2022) utilizes multiple autoencoders (AEs) to model the reverse denoising generation yet only leverages the first AE for interaction prediction during the inference phase. Bénédict et al. (2023) propose 1D binomial diffusion to explicitly model the binary user-level interaction vectors with a Bernoulli process. GiffCF (Zhu et al., 2024) converts the user-level interaction vector into an item-item similarity graph, and employs a smoothed graph-based diffusion-denoising process using the heat equations, which leverages the advantages of both diffusion models and graph signal processing.
@@ 3.2.2. Diffusion for Context-aware Recommendation
First and foremost, a number of works aim to involve temporal/sequential information for diffusion-based context-aware recommendation models (Yang et al., 2024c; Li et al., 2023a; Du et al., 2023; Wang et al., 2024c; Niu et al., 2024; Wang et al., 2024d; Li et al., 2024c). These studies intend to conduct conditional diffusion-based generative recommendation by taking the representations of user behavior sequences as the guidance for reverse denoising process. For example, DreamRec (Yang et al., 2024c) uses a Transformer (Vaswani et al., 2017) encoder to create guidance representations from user behavior sequences and employs a diffusion model to explore the underlying distribution of item space, generating an oracle next-item embedding that aligns with user preferences without the need for negative sampling. As a follow-up work, DiffRIS (Niu et al., 2024) improves the guidance representations from behavior sequences by explicitly modeling the long- and short-term user interests via delicately designed multi-scale CNN and residual LSTM modules. DimeRec (Li et al., 2024c), on the other hand, strives to optimize the diffusion-denoising process. To be specific, for each training sample, the model introduces noise to the target item embedding using geodesic random walk (De Bortoli et al., 2022) on a spherical space, ensuring isotropic and small noise insertion that approximates Gaussian distribution in the tangent space. The reverse denoising process is conducted under the guidance of user sequence representation, and is jointly optimized with the objective of traditional sequential recommender in a multi-task manner. Moreover, instead of integrating the sequential behaviors into one compact representation beforehand for guidance, other works (Du et al., 2023; Li et al., 2023a; Wang et al., 2024c) attempt to directly incorporate a sequence of historical item embeddings during the reverse generation process, resulting in implicit conditions on the sequential information.
@@ 3.2.3. Diffusion for Other Applications
While collaborative and context-aware diffusion-based recommendation models focus on precisely estimating the user preference towards a target item given the in-domain recommendation data (i.e., interaction records and contextual information), diffusion models are also adapted as the core generative functions for other online application tasks. These tasks are closely related to the recommendation scenarios but somehow differ from the aforementioned two categories (i.e., collaborative or context-aware recommendation) in problem formulations and solution paradigms. Hence, we generally categorize them as other applications and briefly introduce these research works as follows.
@@
DMSR (Tomasi et al., 2024) attempts to optimize the slate recommendation with the help of diffusion models. Different from traditional recommendation models that usually estimate the user preference towards each individual item, the slate recommendation aims to optimize the entire collection of items (i.e., a slate/group/bundle) presented to the user at once. Hence, the model should comprehensively consider the in-slate mutual effects (e.g., diversity) and overall utilities, thereby suffering from the complex combinatorial choice space. To this end, diffusion models turn out to be a promising solution to directly generate a set of latent vectors to construct the slate via item mapping. DMSR adopts the diffusion transformer architecture (Peebles and Xie, 2022), and conducts the reverse denoising process over the corrupted in-slate item representations under the guidance of contextual information like user preferences or textual queries. The denoised latent vectors are then decoded back into the discrete item space, resulting in a slate recommendation that maximizes user satisfaction by balancing the relevance and diversity.
@@ 3.2.4. Discussion
Due to the exceptional generative capabilities and flexibility with different backbones and downstream tasks, diffusion-based recommendation models have turned out one of the the state-of-the-art generative recommendation paradigms by learning the underlying distribution of user preferences and item characteristics. From the aforementioned discussion, we can clearly observe the potential advantages through the adaptation of diffusion models as recommenders, including but not limited to capturing the complex and non-linear relationships between users and items and therefore generating diverse and novel recommendations. In this section, as shown in Figure 5, we would like to further discuss this line of research from the following three key perspectives when constructing the diffusion-based recommendation models: (1) what to diffuse, (2) what is the guidance, and (3) how to accelerate.
@@
What to diffuse? As shown in Figure 5, there are generally two types of inputs for the diffusion process: discrete user preference distribution and continuous latent representations. The first choice is to diffuse the interaction matrix or the probabilistic distribution of user preferences over the entire item space (Bénédict et al., 2023; Wang et al., 2023b), e.g., collaborative signals $\mathbf{x}_{0}=\{0,1\}^{|\mathcal{I}|}$. Although such kind of methods is simple and straightforward to directly denoise and generate the target user preference distribution towards potentially unobserved interactions, it is not scalable when facing a large volume of items due to the tremendously large input space and extensive resource requirements. Moreover, other contextual information like multi-modal knowledge can only be involved in the guidance part, which limits the model design. As a result, the second choice proves a more coherent, scalable and commonly adopted solution to conduct the diffusion-denoising process in the continuous latent space (Yu et al., 2023a; Wang et al., 2024e; Li et al., 2024f; He et al., 2024a). In this way, the diffusion-based recommenders are able to first integrate multi-source knowledge (e.g., spatial-temporal data and social networks in addition to collaborative signals) into a unified representation via specially designed encoders, and employ the diffusion model to learn the underlying latent distributions, which is more compatible with various tasks like POI recommendation or social recommendation. The denoised latent representations, usually serving as the user preference vectors or target item representations, are then used to retrieve the top-$K$ relevant items through vector matching for the final ranked item list.
@@
What is the guidance (optional)? As depicted in Figure 5, the guidance serves as the conditioning information to tailor the reverse denoising process for either discrete user preference distributions or continuous latent representations. It plays a crucial role in ensuring that the generated recommendations are not only high-quality but also relevant and personalized for each user. Although optional, the guidance mechanism has been widely adopted in diffusion-based recommendation models, especially for context-aware recommendation where various context information can be fused into the conditioning signals, e.g., sequential information (Yang et al., 2024c; Li et al., 2024c), multi-modal knowledge (Yu et al., 2023a; Wang et al., 2024e), and spatial-temporal data (Qin et al., 2023; Long et al., 2024; Wang et al., 2024e).
@@
How to accelerate? By determining the aforementioned two key factors (i.e., contents to be diffused and the optional guidance), one can already construct the overall architecture of diffusion-based recommendation models. However, the inference efficiency is another important topic when directly adapting diffusion models as recommenders, since online recommender systems are usually real-time services and extremely time-sensitive, where the user request should be responded within around tens of milliseconds (Lin et al., 2024b, 2023a). Diffusion models can be computationally expensive due to the multi-step denoising process, especially when dealing with large-scale recommender systems. Several strategies have been proposed to balance the trade-off between recommendation quality and computational efficiency. For example, Lin et al. (2024a) and Wang et al. (2023b) propose to start the denoising inference from a meaningful input (e.g., outputs from the previous stages or partially corrupted user preference representations) instead of the pure Gaussian noise in traditional diffusion models, which does not only avoid totally eliminating the important personalized information, but also accelerate the inference phase with much fewer denoising steps. Other works also attempt to speed up the denoising process by adopting the early-stopping strategy (Lin et al., 2024a) or one-step inference (Yu et al., 2023a). Despite these preliminary explorations, there is still a lack of systematic studies on the general acceleration strategies for the inference of diffusion-based recommendation models, formulating a promising direction for future research over which we would cast further discussions in Section 4.
@@ 3.3. Diffusion for Content Presentation
The two stages discussed above (i.e., data engineering & encoding, and recommendation models) concentrate on selecting and arranging the optimal ranked item list for recommendation. In this section, the phase of content presentation focuses on customizing and seeking the best presentation strategy of recommended items for different users (e.g., individualized titles or thumbnails). This is non-trivial to not only consider the user’s potential preferences based the historical and contextual information but also involve the visual appearance correlation and coherence. While earlier attempts for content presentation heavily rely on the manual design and pre-defined page arrangement, the emerging of top-tier generative models, especially diffusion models like Stable Diffusion (Rombach et al., 2022b), points out a promising direction to automate and even personalize the content generation for the item display. According to the different types of contents to be generated by diffusion models, the existing research works in this line roughly exhibits three progressive development stages: (1) fashion generation, (2) ad creative generation, and (3) general content generation.
@@ 3.3.2. Ad Creative Generation
Compared with fashion outfits, ad creative generation covers a broader range of visual contents, including but not limited to apps, games, cosmetics, and other promotional images. Previously, the creatives of ad campaigns are designed with great care and manual efforts. However, this human-designer-centered approach are non-scalable and suboptimal, because the designer often struggles to fully consider the global preferences of targeted user groups, and also fails to adapt to different online advertising scenarios. A common solution is to show the original product image with additional design elements adaptively generated by some neural models, e.g., language models to customize the headlines and captions (Mita et al., 2023; Kanungo et al., 2021). Moreover, since the ad creative aims to promote the sale or click rate of a specific product, there are usually more constraints for automatic ad creative generation. For example, we have to maintain the original target product within the generated or edited creative image. Besides, the seller and designer might have additional guidelines and instructions when employing artificial intelligence generated content (AIGC) for creative modification or generation, e.g., the overall visual styles and color themes should be coherent and consistent. To this end, diffusion models are becoming increasingly popular to facilitate the ad creative generation with inpainting/outpainting mode (Lugmayr et al., 2022) and text-to-image controllable generation (Abdollahpouri et al., 2017).
Shilova et al. (2023) propose the generative creative optimization task, and utilize user interest signals to personalize the ad creative generation with the outpainting mode of stable diffusion (Rombach et al., 2022b). The proposed AdBooster first masks out the background of the original product image (e.g., a model with certain fashion outfits), and then finetunes a stable diffusion model to outpaint the masked background with textual prompt guidance based on the user query and contextual information.
@@
More recently, Czapp et al. (2024) employs diffusion models for the creation of eye-catching personalized product images to increase user engagement with recommendations in online retargeting campaigns for real-world industrial applications. Although still adopting the basic fill-in-the-background paradigm for ad creative generation, the authors improve the overall pipeline from the following aspects. Firstly, the position and size of the masked product object would be further adjusted to the center of the image according to the placement for emphasis. Secondly, an optional edge detection is introduced to further reinforce the contours with the background mask. Finally, the authors use LinUCB (Li et al., 2010; Chu et al., 2011) contextual multi-armed bandit algorithm to select the prompt with the highest predicted CTR from a pre-defined pool of prompts associated with the product category in the given context (defined by user, item, and location features), which is designed to balance the generative personalization and online latency constraint.
@@ 3.3.3. General Content Generation
The exploratory research works (Mukande et al., 2024; Shen et al., 2024; Wang et al., 2023a) generally follow the similar framework, where the diffusion-based content generator is guided by the conditions extracted by a hybrid instructor module. The content to be generated by diffusion models is no longer limited to images of clothing or ad creatives, but has already extended to a broader range of modalities. The hybrid instructor module usually involves large language models (LLMs) to deepen the understanding of user intents assisted by other traditional models like sequence models and graph-enhanced models. Built upon this basic setting, Wang et al. (2023a) propose GereRec to further refine the formulation and deepen the integration of the diffusion-based AIGC module with the overall recommendation pipeline. GereRec not only generates the personalized content, but also supports a variety of operations to further adjust the item content to maximize the user-oriented utilities, e.g., thumbnail generation and selection, caption customization, and domain-specific fidelity checks. Although the research in this line is still preliminary and not well validated on large-scale platforms, we believe this is a direction worth delving into, which could potentially lead to the next-generation generative recommender paradigms.